What volume would 1.90 × 1023 molecules of an ideal gas occupy at STP?
1
Expert's answer
2018-12-09T03:57:10-0500

Answer:

n = N/Na = 1.90x10^23 / 6.02x10^23 = 0.306 mole

V = n*Vm

V = 0.306*22.4 = 6.86 litre